OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, 2) with slope 3

OpenStudy (anonymous):

HI WELCOME TO OPENSTUDY :)) slope intercept form is \(\Large y=mx+b\) where m is your slope and b is your y-intercept. put you slope, m=3 into the equation \(\Large y=3x+b\) now, using the point you have, solve for b can you do that?

OpenStudy (anonymous):

I dont know how and i need help

OpenStudy (anonymous):

(4, 2) = (x, y) x=4 y=2 substitute these values into y=3x+b then tell me what you have and i'll tell you what you need to do next?

OpenStudy (anonymous):

okay im working on the problem now(: give me a minute

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

I got 2 = 12b

OpenStudy (anonymous):

i think i did it wrong

OpenStudy (anonymous):

almost right it should be \(\Large 2=12 + b\) now, to solve for b, subtract both sides by 12, what will you have?

OpenStudy (anonymous):

b = -10

OpenStudy (anonymous):

right! (^_^) so you m=3 , b= -10 put this in the equation \(y=mx+b\) you'll get?

OpenStudy (anonymous):

Almost but the options that i have for this problem are a. y-2=3(x+4) b. y-2=3(x-4) c.y-2=-3(x-4) d.y+2=3(x+4) and I dont understand how they got the choices

OpenStudy (anonymous):

aah, they use this : \((y-y_1)=m(x-x_1)\) you have \((x_1, y_1)=(4, 2)\) and \(m=3\) substituting these: you will have \(\LARGE y-2=3(x-4)\) if you will simplify this, you'll get the answer you have before, but i see that they did not simplify it, so you can have this as your answer ;)

OpenStudy (anonymous):

do you have time to help me with 7 more questions ? i have an online class and i only have 29 more minutes to finish them and i really could use the help.

OpenStudy (anonymous):

oh, i'm so sorry, i'm afraid i can't, i have classes tomorrow and i have to sleep now. just post your questions again and maybe there are other users that can help you :)

OpenStudy (anonymous):

okay thanks for your help :)

OpenStudy (anonymous):

no problem :) goodnight